1. Specialization and Trade Alberto, Bob, and Sichao each produce xylophones and yarn. The daily production possibilities frontiers for each of them are given in the following graphs Albertos daily PPF Bobs daily PPF Sichaos daily PPF 4 4 4 0 1 2 3 4 5 0 1 2 3 4 5 0 1 2 3 4 5 xylophones ylophones xylophone Circle one.) a. Who has the absolute advantage in the production of xylophones? Alberto Bob Sichao b. Who has the absolute advantage in the production of yarn? c. Fill in the table with the appropriate opportunity costs for each of the three producers Alberto Bob Sichao Circle one.) opportunity cost of 1 xylophone opportunity cost of 1 unit of yarm Alberto Bob Sichao d. For each potential pair of trading partners below, circle the one who would produce xvlophones if those two were to completely specialize and trade between each other (without the third person). In each case, briefly explain the producer you circled would specialize in the production of xylophones Alberto and Bob ii) Alberto and Sichao ili Bob and Sichao
Answer #1

(a) Sichao has an absolute advantage in the production of xylophones because Sichao produces more xylophones ( i.e 4 )with the same resources than Alberto and Bob.

(b) Bob has an absolute advantage in the production of yarn because Bob produces more yarn (i.e 5 ) with the given resources than Alberto and Sichao.

(c)

Opportunity cost of 1 xylophone Opportunity cost of 1 units of yarn
Alberto 1/3 = 0.33 3
Bob 5/2= 2.5 2/5 =0.4
Sichao 4/4=1 4/4=1

(d) (i) For the pair Alberto and Bob : Alberto has a comparative advantage in the production of xylophone because he has the lower opportunity cost for the production of xylophone than Bob.Therefore, Alberto would completely specialize in xylophone.

(ii) For the pair Alberto and Sichao : Alberto has a comparative advantage in the production of xylophone because he has the lower opportunity cost for the production of xylophone than Sichao.Therefore, Alberto would completely specialize in xylophone.

(iii) For the pair Bob and Sichao: Sichao has a comparative advantage in the production of xylophone because he has the lower opportunity cost for the production of xylophone than Bob.Therefore, Sichao would completely specialize in xylophone.